OpenStudy (anonymous):

4^(x + 7) = 6^(x – 1)

OpenStudy (callisto):

4^(x + 7) = 6^(x – 1) (x+7) log4 = (x-1) log6 xlog4 + 7log4 = xlog6 -log6 x( log4 -log6) = - (log6 + 7log4) x = - (log6 + 7log4) / ( log4 -log6) =?